Grade 11Mechanics

Let a=2;b=3;c=6.Angle between any two vectors is 120 and the vectors are coplaner,find |ā+b+c|.

Tushar Bailwal
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
(a + b + c) . ( a + b+c) = a^2+ b^2 + c^2 + 2a.b + 2b.c + 2c.a
 
|a + b + c|^2 = 4 + 9 + 36 – 6 – 18 – 12 = 13
 
hence |a + b + c| = sqrt (13)